OpenStudy (anonymous):

Find the product of the following complex numbers x=3(cosπ/4 + sinπ/4) y=7(cos3π/8 + sin 3π/8) Express your answer in trigonometric form.

OpenStudy (jdoe0001):

same as before then \(\bf \cfrac{r_1[cos(\alpha)+isin(\alpha)]}{r_2[cos(\beta)+isin(\beta)]}\implies \cfrac{r_1}{r_2}[cos(\alpha - \beta)+isin(\alpha - \beta)]\)

OpenStudy (jdoe0001):

h... shoot is the product... well ...

OpenStudy (jdoe0001):

\(\bf r_1[cos(\alpha)+isin(\alpha)]\cdot r_2[cos(\beta)+isin(\beta)] \\ \quad \\ \implies r_1\cdot r_2[cos(\alpha + \beta)+isin(\alpha + \beta)]\)

OpenStudy (anonymous):

There we are. Thanks. I just couldn't find that equation anywhere!

OpenStudy (jdoe0001):

yw